# Assortment Optimization for Patient-Provider Matching

### v1

@@ 6 Conclusion and Discussion
The patient-provider relationship is key to quality healthcare, yet provider turnover frequently leads to patients being without providers. Our work studies how to re-match patients in such a scenario through assortment optimization. Our study proposes algorithmic approaches for patient-provider matching systems and offers analytical and empirical insights to inform system design. We further offer contributions to one-shot matching under random response order, an understudied phenomenon within the assortment literature. We conclude by providing recommendations for deploying assortment-based systems into practice and discussing extensions of our work.
Recommendations for Deployment
@@
    Larger assortments become increasingly useful as patients become more selective because patients are more likely to reject matches, making it important to give more options. Moreover, assortment sizes should be tailored to patient behavior; the choice model and match probability impact which policy performs best.
@@
    The level of heterogeneity impacts the amount of work needed to design a good algorithm. For example, patients with diverse preferences require less work to find the ideal match, and even relatively simple baselines would work well. However, homogeneous scenarios (e.g., a small city with few providers) require care when matching patients and providers. For example, when patients all prefer the same provider, simple solutions such as the pairwise policy can perform poorly.
@@
Model Extensions
We detail several extensions of this model that better capture real-world patient-provider matching. As noted in Section 2.1, we assume that the response time for each patient is i.i.d. A natural extension would be to develop policies that work well independent of the response order distribution. Our experiments also assume that the choice model, $f_{i}$, is constant across patients. However, in reality, patients might have heterogeneous choice models. For example, in the uniform choice model, the value of $p$ could vary across patients. It would be interesting to see how extensions of our policies would perform under heterogeneous choice models.
Extensions to Other Domains
Future work can extend our framework to other scenarios with random response orders. Examples include matching between students and courses, where universities could offer each student an assortment of courses, and between schools and children, where school choice might lead counties to offer each child multiple options (Chen and Sönmez, 2006).
@@ Acknowledgements
We thank Aakash Lahoti, Keegan Harris, and George Chen for their comments. We thank the reviewers at ML4H for their feedback on a preliminary version of this work. We thank Susan Barrett and Karen Hall for their input from a health system perspective. Co-author Raman is supported by an NSF GRFP Fellowship.
Appendix A Extended Experiments
More patients than providers
Figure 9: We compare the match rate across policies when varying the patient/provider ratio, and find that the greedy and gradient descent policies perform best, while the group-based and pairwise policies lag.
We compare the match rate across policies from Example 5 and find that the gradient descent and greedy policies achieve the highest match rate for large patient/provider ratios (see Figure 9). As shown in Section 3.1, the greedy policy achieves the highest match rate for the uniform choice model, so it is unsurprising that it achieves the highest match rate. Beyond that, the high performance of gradient descent matches the trends with match quality in Example 5.
Impact of $p$ and $\theta$
Figure 10: For $p\leq 0.5$, we find that all four policies achieve the same match rate, while for $p\geq 0.75$, we find that gradient descent achieves a slightly lower match rate compared to the greedy, pairwise, and group-based policies.
We extend Example 6 to compare the match rates across policies when varying $p$ and $\theta$ in Figure 10. All non-random policies have similar match rates for $p\leq 0.5$, implying that match quality is better at differentiating policies than match rate. For larger $p$, the gradient descent policy performs slightly worse in match rate, while the greedy, pairwise, and group-based policies achieve the same match rate. As in Example 6, we find that the best-performing policies for match quality also tend to exhibit the highest match rate (in this case, the group-based policy is optimal for both).
Misspecified Choice Models
@@ Appendix A Extended Experiments
Figure 11: We evaluate the performance of policies when the observed match probability, $\hat{p}$, differs from the true match probability, $p$. We find that the observed match probability has little impact on the relative ordering of policies, although the group-based policy improves as $p$ and $\hat{p}$ get closer.
@@
Our experiments in Section 5 consider situations with known values of $p$, but real-world scenarios frequently involve misspecified choice models. To model this, we run policies with some observed match probability $\hat{p}$ while letting $p$ be the match probability. This setup allows us to simulate real-world scenarios where match probability is unknown. We compare policies across $\hat{p}\in\{0.1,0.25,0.5,0.75,0.9\}$ and $p\in\{0.1,0.25,0.5,0.75,0.9\}$.
In Figure 11, we demonstrate that most policies perform similarly across $\hat{p}$ for fixed $p$. This trend holds because the random, greedy, and pairwise policies are unimpacted by $\hat{p}$. We find that the group-based policy gets worse as $\hat{p}$ and $p$ differ; for example, when $p=0.9$ and $\hat{p}=0.1$, the group-based policy performs worse than the pairwise policy, while when $\hat{p}=0.9$, the policies perform similarly. We note that the differences are slight even for the group-based policy, so our policies are generally robust to misspecification in match probabilities.
Appendix B Proofs
See 2.3
@@ Appendix B Proofs
Suppose that $s$ patients are offered the provider $j=1$: $\lVert\mathbf{X}\rVert_{1}=s$. In the uniform choice model, the probability that any of the $s$ patients select the single provider is $p$, so the probability of provider $j=1$ matching is $1-(1-p)^{s}$. By symmetry in the response order, the chance that each patient is selected is equal, and so the match quality is
@@
See 3.1
@@
$p(\Delta(N-1)+1)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (21)
Next, consider the greedy policy where $\pi^{R}(\theta)=\mathbf{1}$. All patients prefer provider $1$, and because of this, all patients have an equal likelihood of matching with provider $1$. Each patient has a $\frac{1}{N}$ chance of matching with provider $1$, and therefore the expected match quality for provider $1$ is $\frac{1}{N}+\frac{N-1}{N}(2\Delta)$. For all the $N-1$ other providers, we receive a reward of $\Delta$ upon matching, and each patient matches with probability $p$; therefore, we receive a total match quality of at most $\frac{1}{N}+2\frac{N-1}{N}\Delta+p(N-1)\Delta$. Taking the ratio of the greedy and optimal policies yields the following:
$\frac{\frac{1}{N}+\frac{2\Delta(N-1)}{N}+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p}$
$\leq\frac{1}{Np}+\frac{2\Delta}{p}+N\Delta$
Finally, letting $N=\frac{3}{\epsilon p}$ and $\Delta=\min(\frac{\epsilon}{3N},\frac{p\epsilon}{6})$ yields that
$\frac{1}{Np}+2\frac{\Delta}{p}+N\Delta\leq\frac{\epsilon}{3}+\frac{\epsilon}{3}+\frac{\epsilon}{3}=\epsilon$ (22)
Therefore, for any choice of $\epsilon$ and $p$, there exists a choice of $\theta$ so the greedy policy is at most an $\epsilon$ approximation. ∎
See 9
@@
Next, we upper bound the performance of the optimal policy. The optimal policy achieves a match quality of
For any order $\sigma$, we have that no provider can be selected by two patients, and each patient can select at most one provider. That is, for any ordering, we can bound the match quality as
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (24)

### v2

@@ 6 Conclusion and Discussion
The patient-provider relationship is foundational for quality healthcare, yet provider turnover frequently leads to patients being without providers. Our work studies how to re-match patients in such a scenario by reformulating the problem through assortment optimization. Our study proposes algorithmic approaches for patient-provider matching systems and offers analytical and empirical insights to inform system design. Moreover, these policies broadly apply to one-shot matching under random response order, an understudied phenomenon within the assortment literature. We conclude by providing recommendations for deploying assortment-based systems into practice and discussing extensions of our work.
6.1 Managerial Insights
@@
    Larger assortments become increasingly useful as patients become more selective because patients are more likely to reject matches, making it important to give more options. Moreover, assortment sizes should tailor to patient behavior; the choice model and match probability impacts the best policy.
@@
    Various metrics are in tension when constructing assortments. In many situations, we cannot simultaneously maximize match rate and match quality, and incorporating fairness only complicates the picture. Ultimately, these models provide a slate of options for practitioners to consider when weighing the priority of various metrics.
@@
6.2 Broader Impact and Extensions
Our work proposes policies for assortment optimization in the one-shot random response setting, and future work can extend these policies to other models within patient-provider matching and domains. Within patient-provider matching, natural extensions include incorporating other metrics, including provider workload and total response times. Other extensions include modeling a dynamic pool of providers, where providers might enter and exit throughout the matching process and incorporating side constraints for individual patients or providers.
Outside of patient-provider matching, our policies apply to a diverse class of problems with offline assortments and online random responses. Such a setting is a feature of many two-sided markets, and we present two examples. The first is assortment allocation in food delivery platforms; these platforms offer trips that drivers pick up and respond to in some random order, and batch offering eases the central planner’s logistical burden. A second example is the matching between students and courses, where universities could offer each student an assortment of courses, and between schools and children, where school choice might lead counties to offer each child multiple options (Chen and Sönmez, 2006). These examples highlight the broader relevance of our problem setting and the applicability of our policies across these domains.
@@ Acknowledgements
We thank Aakash Lahoti, Keegan Harris, Kevin Li, Ariana Tang, Fei Fang, and George Chen for their comments. We thank the reviewers at ML4H for their feedback on a preliminary version of this work. We thank Susan Barrett and Karen Hall for their input from a health system perspective. Co-author Raman is supported by an NSF GRFP Fellowship.
Appendix A Misspecification Experiments
Misspecification in $p$
@@ Appendix A Misspecification Experiments
Figure 11: We evaluate the performance of policies when the observed match probability, $\hat{p}$, differs from the true match probability, $p$. We find that the observed match probability has little impact on the relative ordering of policies.
@@
Our experiments in Section 4 consider situations with known values of $p$, but real-world scenarios frequently involve misspecified choice models. To model this, we run policies with some observed match probability $\hat{p}$ while letting $p$ be the match probability. This setup allows us to simulate real-world scenarios where match probability is unknown. We compare policies across $\hat{p}\in\{0.1,0.25,0.5,0.75,0.9\}$ and $p\in\{0.1,0.25,0.5,0.75,0.9\}$ with $M=25$ and $N=100$. In Figure 11, we demonstrate that policies perform similarly across $\hat{p}$ for fixed $p$. This trend holds because the random, greedy, and pairwise policies are unimpacted by $\hat{p}$, while other policies are not sensitive to the exact choice of the match probability. These results imply our policy’s robustness to their exact knowledge of the choice model.
Variance in $p$ and $\theta$
Figure 12: We compare the performance of policies when either $\theta$ or $p$ is misspecified by some noise $\epsilon$. We find no noticeable differences between policy performances even as noise size is increased; that is, the relative ordering of policies remains constant.
We consider the impact that noisy $p$ and $\theta$ have on the performance of policies. For each parameter, we modify it, so $p_{i}=p+\epsilon$ for some unknown noise $\epsilon$, and similarly, we let $\theta=\hat{\theta}+\epsilon$, where we only observe $\hat{\theta}$. We let $\epsilon=\mathcal{N}(0,s^{2})$ for $s\in\{0,0.05,0.1,0.2\}$, and similarly for $\epsilon$, we let $\epsilon=\mathcal{N}(0,s^{2})$ for $s\in\{0,0.05,0.1,0.2\}$. We again let $M=25$ and $N=100$.
In Figure 12, we demonstrate that policies perform similarly across values of $s$. For both $p$ and $\theta$, we find that the policies that perform well under the noiseless scenario similarly perform well in the noisy scenario. For example, in all scenarios, we find that the gradient descent policy performs best, while the pairwise policy performs worst. This indicates that our results are not very sensitive to the exact knowledge of $p$ and $\theta$.
Appendix B Policies with Non-Unit Capacities
We develop variants of each policy that work under non-unit capacity $c_{j}$. For greedy, we leave the policy as is, as it is not dependent on the unit capacity assumption. For both the pairwise and group-based assumption, we relax the initial linear programming solution so it has $\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\leq c_{j}$, but proceeds as normal otherwise. Finally, for the gradient descent policy, we modify the computation of $h(\mathbf{X})_{i,j}$. Recall that $h(\mathbf{X})_{i,j}$ is an upper bound on the availability of provider $j$ for patient $i$. Originally, at position $t$, the availability probability is $(1-p^{\prime}_{j})^{t-1}$, which represents the probability that none of the first $t-1$ patients select provider $j$. Here, we now modify this so that no more than $c_{j}-1$ patients select provider $j$ in the first $t-1$ patients. We compute this by approximating the number of patients who select provider $j$ through a normal approximation $\mathcal{N}(p^{\prime}_{j}(t-1),p^{\prime}_{j}(1-p^{\prime}_{j})(t-1))$. Then we want to compute the probability that $\mathbf{Pr}[\mathcal{N}((t-1)p^{\prime},(t-1)p^{\prime}(1-p^{\prime}))\leq c_{j}-1]$. This is equivalent to $\Phi(\frac{c_{j}-1-(t-1)p^{\prime}}{\sqrt{(t-1)p^{\prime}(1-p^{\prime})}})$, which we then lower bound as:
$\left\{\begin{array}[]{ll}\displaystyle\frac{z}{z^{2}+1}\cdot\frac{1}{\sqrt{2\pi}}e^{-z^{2}/2}&\text{if }c_{j}-1\leq(t-1)p^{\prime}\\ \\ (1-p^{\prime})^{t-1}&\text{otherwise}\end{array}\right.$ (26)
Here, $z=\frac{c_{j}-1-(t-1)p^{\prime}}{\sqrt{(t-1)p^{\prime}(1-p^{\prime})}}$, and we use a bound on $\Phi(z)$ (Cook, 2009).
Appendix C Varying Provider Capacity
Figure 13: We vary the capacity for providers from $1$ to $4$, and find that increased provider capacity leads to homogenization of policy performance. We find that the same effect can be accomplished when expanding the number of providers as well; having 25 providers each with a capacity of 4 yields similar results to having 100 providers.
Following the results in Section 4.3, we run experiments to understand the impact of provider capacity on performance. We vary provider capacity from 1 to 4 (so each provider can take on 1 to 4 patients) when fixing $N=100$ and $M=25$, and plot this in Figure 13. We find that as the amount of provider capacity is increased, policy performances are increasingly homogeneous. As providers have a larger capacity, individual assortments matter less, as all patients have a higher chance of matching with their most preferred provider. Moreover, in Figure 13 bottom, we find that increasing provider capacity produces a similar result to increasing the total number of providers. For example, having $M=25$ providers each with a capacity of 4 produces similar results to having $M=100$ providers.
Appendix D Batching Algorithm
We study how to construct batches so to maximize the match quality. Formally, suppose we construct $L$ batches, $b_{1},b_{2},\ldots,b_{L}$, where $b_{k}\subseteq[N]$, and we make offers first to batch $b_{1}$. Each set of batches modifies the response order distribution so that $\sigma$ is now distributed according to $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$, where $S$ is some probability distribution function. Throughout this section, let $\mathrm{MQ}(\mathbf{X},\theta,f,\sigma)=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\mathbf{X}\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$. If we can find some optimal ordering $\sigma^{*}=\operatorname*{arg\,max}_{\sigma}\mathrm{MQ}(\mathbf{X},\theta,f,\sigma)$, then our goal is to set $b_{1},b_{2},\ldots,b_{L}$ so $\sigma^{*}$ and $S(b_{1},b_{2},\ldots,b_{L})$ are similar.
Finding the optimal ordering $\sigma^{*}$
To understand $\sigma^{*}$, we first introduce a policy $\pi^{A}(\theta,\sigma)$, which improves the pairwise policy with knowledge of the order $\sigma$. Recall that $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$, and let $v^{-1}_{j}=i$ if $v(\theta)_{i}=j$ and $\sigma^{-1}_{i}=t$ if $\sigma_{t}=i$. We can compute the order $\sigma^{*}$ that optimizes $\mathrm{MQ}(\pi^{A}(\theta),\theta)$. $\pi^{A}$ adds to the pairwise assortment based on knowledge of the response order $\sigma$. For example, if $\sigma^{-1}_{i}=1$ and $\sigma^{-1}_{i^{\prime}}=2^{\prime}$, then we know that offering $v(\theta)_{i}$ to provider $i^{\prime}$ can only improve the match quality because $v(\theta)_{i}$ will only be available if unselected by patient $i$. More generally, we add on pairs $(i^{\prime},v(\theta)_{i})$ based on the order $\sigma$ when $\sigma^{-1}_{i}<\sigma^{-1}_{i^{\prime}}$. Intuitively, if patient $i$ comes earlier than patient $i^{\prime}$ in $\sigma$, then it’s safe to offer $v(\theta)_{i}$ to patient $i^{\prime}$ as well because patient $i$ will have the first opportunity to claim $v(\theta)_{i}$. More formally, we include $(i^{\prime},v(\theta)_{i})$ in the augmented assortment if: a) patient $i^{\prime}$ prefers $v(\theta)_{i}$ at least as much as their original provider under $\pi^{P}$, i.e., $\theta_{i^{\prime},v_{i}}\geq\theta_{i^{\prime},v_{i^{\prime}}}$, and b) $i$ comes before $i^{\prime}$ in $\sigma$, i.e., $\sigma^{-1}i<\sigma^{-1}{i^{\prime}}$. We focus on such a policy because it represents a natural extension of the pairwise policy to incorporate the order $\sigma$. We next characterize $\sigma^{*}$ for the policy $\pi^{A}$ then use this to construct the batches $b_{k}$:
Lemma D.1.
Let $G=(V,E)$ be a graph with $N$ nodes such that there is an edge from $i$ to $i^{\prime}$ if $\theta_{i,v(\theta)_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v(\theta)_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Let $\sigma^{*}$ be the optimal ordering for the policy $\pi^{A}$. Then, traversing the nodes defined by $\sigma^{*}$ is a reverse topological ordering on $G$.
Intuitively, the graph $G$ captures who envies whom: if there’s an edge from $i$ to $i^{\prime}$, then $i$ would prefer the provider assigned to $i^{\prime}$, suggesting $i^{\prime}$ should come earlier in the response to minimize blocking. We additionally let $v(\theta)_{i}=-1$ for unmatched patients. We demonstrate this through casework on pairs of nodes $i,i^{\prime}$ in $\sigma^{*}$.
Constructing batches $b$
We next construct $b_{1},b_{2},\ldots,b_{L}$, so $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ is close to the optimal ordering $\sigma^{*}$ for $\pi^{A}$. The idea is to partition $\sigma^{*}$ into $L$ batches; then differences only emerge within batches rather than between batches. We note that while we can approximate $\sigma^{*}$ through topological ordering, computing the optimal $\sigma^{*}$ is not guaranteed (as there can be many topological orderings). Once we have an approximation of $\sigma^{*}$, we can use dynamic programming to find the set of batches $L$ that minimizes the number of edges within a batch. By minimizing the number of edges within a batch, we reduce stochasticity in $\sigma$, as edges between batches have a fixed order. Formally, we can guarantee that, given knowledge of $\sigma^{*}$, we can approximate this through batching for sufficiently large $L$ under a set of regularity conditions:
Theorem D.1.
Let $\sigma^{*}$ be an optimal ordering. Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v(\theta)_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v(\theta)_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Suppose that there exists a partition of $\sigma^{*}$ into $L$ batches, $b_{1},b_{2},\ldots,b_{L}$, such that for any partition $b_{k}$, no $i,i^{\prime}\in b_{k}$ have a common descendant in $G=(V,E)$. Then $\pi^{A}$ achieves the same match quality under partition $g$ as under the optimal ordering:
$\mathbb{E}_{\sigma\sim S(b_{1},b_{2},\ldots,b_{L})}[\mathrm{MQ}(\pi^{A}(\theta),\theta,\sigma)]=\mathrm{MQ}[\pi^{A}(\theta),\theta,\sigma^{*}]$ (27)
We prove this by demonstrating that, if we can control for the descendants of each patient, then we can fully control the order $\sigma$, and let it be equal to the optimal order $\sigma^{*}$. Here, the common descendant condition ensures that the only source of uncertainty for the ordering $\sigma$ is the order of the batches themselves rather than any uncertainty for ordering within a batch. Overall, the takeaway is that in some situations, controlling the batches $b$ is equivalent to controlling the full order $\sigma^{*}$.
Appendix E Semi-Synthetic Experiment Details
Appendix F Proofs
Theorem 2.1 Let $f$ be the uniform choice model with probability $p$. Suppose $M=1$, and let $u_{1},u_{2},\ldots,u_{N}$ be a permutation of $\{1,\ldots,N\}$ such that $\theta_{u_{1}}\geq\theta_{u_{2}}\cdots\theta_{u_{N}}$. Let $s$ be defined as follows:
$s=\operatorname*{arg\,max}_{s}(1-(1-p)^{s})\frac{\sum_{i=1}^{s}\theta_{u_{i},1}}{s}$ (28)
Then the policy which maximizes match rate is $\pi(\theta)=\mathbf{X}=\mathbf{1}$ while the policy which maximizes match quality is $\pi(\theta)=\mathbf{X},\mathbf{X}_{u_{1},1}=\mathbf{X}_{u_{2},1}\cdots\mathbf{X}_{u_{s},1}=1$, where $\mathbf{X}_{i,1}$ is 0 otherwise.
@@ Appendix F Proofs
Suppose that $s$ patients are offered the provider $j=1$: $\lVert\mathbf{X}\rVert_{1}=s$. In the uniform choice model, the probability that any of the $s$ patients select the single provider is $p$, so the probability of provider $j=1$ matching is $1-(1-p)^{s}$. Then the assortment that maximizes the match rate has $p=N$, and so $\mathbf{X}=\mathbf{1}$.
Next, we find the optimal match quality assortment. By symmetry in the response order, the chance that each patient is selected is equal, and so the match quality is
@@
Theorem 3.1 Let $f$ be the uniform choice model with match probability $p$. Then
$\mathrm{MR}(\pi^{R}(\theta),\theta,f)\geq\min(p,\frac{M}{N})$ (30)
Moreover, there exists no policy $\pi^{\prime}$ such that $\mathrm{MR}(\pi^{\prime}(\theta),\theta,f)>\mathrm{MR}(\pi^{R}(\theta),\theta,f)$
Proof.
For the greedy policy, we note that $\pi^{R}(\theta)_{i}=\mathbf{1}$ for all patients $i$. Therefore, the only situation where a patient is unmatched under the uniform choice model is either a) there are no available providers, or b) they fail to match with probability $1-p$. For the former situation, we note that this only occurs when all $M$ providers match, meaning that the match rate is $\frac{M}{N}$. For the latter situation, we note that if all providers fail to match, then each patient has a $p$ chance of matching. Let $\alpha$ be the probability that all providers are taken. Then our overall match rate is $\mathrm{MR}(\pi^{R}(\theta),\theta,f)=\alpha\frac{M}{N}+(1-\alpha)p\geq\min(p,\frac{M}{N})$. Moreover, no policy can perform better because no policy can expand upon the set of assortments offered to patients. ∎
Theorem 3.2 Let $f$ be the uniform choice model with match probability $p$. Then
$\mathrm{MR}(\pi^{P}(\theta),\theta,f)=p\frac{\min(M,N)}{N}$ (31)
Proof.
In the pairwise policy, there are $\min(M,N)$ pairs of patients and providers that are matched up. Each is successfully matched with probability $p$ in the uniform choice model, yielding the overall match probability to be $p\frac{\min(M,N)}{N}$
Lemma F.1.
Let $\pi$ be a policy that augments the pairwise policy; $\pi(\theta)_{i,j}\geq\pi^{P}(\theta)_{i,j}$ for any $\theta$ for all $i,j$. Let $G=(V,E)$ be a directed graph with $N$ nodes such that nodes $i$ and $i^{\prime}$ are connected if $\pi(\theta)_{i,v(\theta)_{i^{\prime}}}=1$ Here, $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$. Then we have that
$\mathrm{MR}(\pi,\theta,f)=\mathrm{MR}(\pi^{P},\theta,f)$ (32)
if each component in $G$ is a complete digraph.
Proof.
Consider patient $i$ in a component that is a connected component of size $k$. For any $j$ with $\pi(\theta)_{i,j}=1$, there exists $k-1$ other $i^{\prime}$ with $\pi(\theta)_{i^{\prime},j}=1$ because all patients within the complete graph have provider $j$ on their assortment. Therefore, at most $k-1$ other patients can select providers from $\pi(\theta)_{i}$, and therefore $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}$ can decrease by at most $k-1$ before $\sigma_{t}=i$. Moreover, because $\min(M,N)$ patients have a non-empty assortment, in expectation, we have that $\min(M,N)p$ patients match.
Theorem 3.3 Let $f$ be the uniform choice model. Then
$\mathrm{MR}(\pi^{G}(\theta),\theta,f)=\mathrm{MR}(\pi^{P}(\theta),\theta,f)$ (33)
Proof.
We prove that the group-based policy matches the assumptions of Lemma F.1, and therefore, has the same match rate as the pairwise policy. We note that, within a group $\mathbf{q}$, all patients share the same assortment. Moreover, this assortment is the union of assortments of all pairwise assortments for each patient. Therefore, we can view each group as a complete graph among all $i$ with $q_{i}=1$, and so the induced graph $G$ consists of complete digraphs. ∎
Theorem 3.4: Let $f$ be the uniform choice model with match probability $p$. Then for any $p<1$, there exists $\theta$ such that the unique maximizer of $\langle g(h(\mathbf{X})),\theta\rangle$, $\mathbf{X}^{*}$, has a match rate:
$\mathrm{MR}(\mathbf{X}^{*},\theta,f)<\mathrm{MR}(\pi^{P}(\theta),\theta,f)$ (34)
Proof.
We will construct an example where the unique maximizer performs worse than the pairwise policy. Consider $N=M=2$, and let $\theta_{1}=[1,0]$ and $\theta_{2}=[1,\epsilon]$. Here, $\langle g(h(\pi^{P}(\theta))),\theta\rangle=1+\epsilon$. Note that any policy that does not offer all providers to some patients performs worse than $\pi^{P}(\theta)$. Next, we will show that $\mathbf{X}_{1}=[1,0],\mathbf{X}_{2}=[1,1]$ outperforms $\pi^{P}(\theta)$. Here, $h(\mathbf{X})_{2,2}=1$ and $h(\mathbf{X})_{*,1}=1-\frac{p}{2}$. Therefore, the expected match quality is $2-p+\frac{p}{2}\epsilon$. Setting $\epsilon<\frac{1-p}{1-p/2}$ (which can always be done when $p<1$) yields that $\mathbf{X}$ is better; moreover, setting $\mathbf{X}_{1,2}=1$ can only lead to a worse match quality. Under $\mathbf{X}$, the match rate then goes from $p$ to $p-\frac{p^{2}}{4}$, showing that $\mathrm{MR}(\mathbf{X},\theta,f)<\mathrm{MR}(\pi^{P}(\theta),\theta,f)$. ∎
Theorem 3.5 Let $f$ be the uniform choice model with match probability $p$. For any $p$ and $\epsilon$, there exists a $\theta$ such that
$\mathrm{MQ}(\pi^{R}(\theta),\theta,f)\leq\epsilon\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (35)
@@
$p(\Delta(N-1)+1)=\mathrm{MQ}(\pi(\theta),\theta,f)\leq\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (36)
Next, consider the greedy policy where $\pi^{R}(\theta)=\mathbf{1}$. All patients prefer the provider $1$, and because of this, all patients have an equal likelihood of matching with the provider $1$. Each patient has a $\frac{1}{N}$ chance of matching with provider $1$, and therefore the expected match quality for provider $1$ is $\frac{1}{N}+\frac{N-1}{N}(2\Delta)$. For all the $N-1$ other providers, we receive a reward of $\Delta$ upon matching, and each patient matches with probability $p$; therefore, we receive a total match quality of at most $\frac{1}{N}+2\frac{N-1}{N}\Delta+p(N-1)\Delta$. Taking the ratio of the greedy and optimal policies yields the following:
$\small\frac{\frac{1}{N}+\frac{2\Delta(N-1)}{N}+p(N-1)\Delta}{p(\Delta(N-1)+1)}\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p(\Delta(N-1)+1)}$ (37)
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p}\leq\frac{1}{Np}+\frac{2\Delta}{p}+N\Delta$ (38)
Finally, letting $N=\frac{3}{\epsilon p}$ and $\Delta=\min(\frac{\epsilon}{3N},\frac{p\epsilon}{6})$ yields that the approximation ratio is $\epsilon$. ∎
Theorem 3.6 Let $f$ be the uniform choice model with match probability $p$. Then
$\mathrm{MQ}(\pi^{P}(\theta),\theta,f)\geq p\mathrm{MQ}(\pi^{*}(\theta),\theta,f)$ (39)
@@
Next, we upper bound the performance of the optimal policy. The optimal policy achieves a match quality of $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ For any order $\sigma$, we have that no provider can be selected by two patients, and each patient can select at most one provider. That is, for any order, we can bound the match quality as
$\mathrm{MQ}(\pi^{*}(\theta),\theta,f)\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}$ (40)
$=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (41)